OpenStudy (anonymous):

solve log(10-4x)=log(10-3)

OpenStudy (anonymous):

Log(10-3)? or is it Log(10-3x)?

OpenStudy (anonymous):

\[\large \log(10-4x)=\log(10-3)\] \[\large \log(10-4x)-\log(7)=0\] Using your logarithm laws: \[\large \log(a)-\log(b)=\log(\frac{a}{b})\] \[\large \log(10-4x)-\log(7)=\log(\frac{10-4x}{7})\] \[\large \log(\frac{10-4x}{7})=0\] Using logarithm laws: \[\log(a)=0\] Can be written as: \[10^0=a\] \[a=1\] Therefore applying this law to your question: \[\large \frac{10-4x}{7}=10^0\] \[\large \frac{10-4x}{7}=1\] \[\large 10-4x=7\] \[\large 4x=3\] \[\large x=\frac{3}{4}\]

OpenStudy (anonymous):

Hi thank you so much! and yeah it was log(10-3x)

OpenStudy (anonymous):

Then that's the not the correct answer. Because I assumed you typed it correctly. Don't copy what I wrote!! You should really check what you write on the internet to see whether you made a typo of some sort. Then people would be able to show you how to solve your problem without giving you an incorrect result.

OpenStudy (anonymous):

Here's the correct worked out solution: \[\log(10-4x)=\log(10-3x)\] \[\log(10-4x)-\log(10-3x)=0\] \[\log(\frac{10-4x}{10-3x})=0\] \[\frac{10-4x}{10-3x}=10^0\] \[\frac{10-4x}{10-3x}=1\] \[10-4x=10-3x\] \[0=x\] Therefore: \[\huge x=0\]